OpenStudy (anonymous):

dy/dx= (3y-2x)/(8y-3x). Find d^2y/dx^2.

OpenStudy (anonymous):

I understand that I am finding the second derivative, and I was wondering if I am able to derive it as 3y/8y - 2x/3x.

zepdrix (zepdrix):

Noooo :O Can't split up denominators like that! You could split it up like:\[\Large \frac{3y}{8y-3x}-\frac{2x}{8y-3x}\]But I don't think that's really going to help us :)

zepdrix (zepdrix):

A fraction is sort of like a house :D The numerator is all the furniture and stuff. We can move the furniture around, but we can't separate the foundation. Ok maybe that was a silly explanation lol

zepdrix (zepdrix):

Looks like we need to apply the `Quotient Rule` for this one, yes?\[\Large y''\quad=\quad \frac{\color{royalblue}{(3y-2x)'}(8y-3x)-(3y-2x)\color{royalblue}{(8y-3x)'}}{(8y-3x)^2}\]

OpenStudy (anonymous):

Would the derivative of (3y-2x) be (3dy/dx-2)? And similar fashion for (8y-3x)' ?

zepdrix (zepdrix):

yes, very good.

zepdrix (zepdrix):

They're probably going to want the final answer in terms of x and y, NOT in terms of x, y and y'. So you'll have a bit of work to do after you find the derivative :( Plugging in the y's from your original problem.

OpenStudy (anonymous):

Ok, just a quick quesiton. When I plugged this into an online calculator *shame shame* it gave me that (3y-2x)'=-2. Is that true?

zepdrix (zepdrix):

If you typed it into wolfram, lemme check really quick. *checking*... No it is not true. Do you see the weird curly d's on the derivative operator? They took a `partial` derivative instead of a normal derivative with respect to x. It's something different that you'll learn about later. Doesn't really apply here though :)

OpenStudy (anonymous):

Ok, so after multiplying everything throughout the numerator, I now need to plug in my original dy/dx into the d^2y/dx^2?

zepdrix (zepdrix):

\[\Large y''\quad=\quad \frac{\color{orangered}{(3y'-2)}(8y-3x)-(3y-2x)\color{orangered}{(8y'-3)}}{(8y-3x)^2}\]Yah you'll need to plug your original setup into each of those `y'`s :o

zepdrix (zepdrix):

You could plug before multiplying everything out, or after, it's up to you :D

OpenStudy (anonymous):

Ok. So I have the point (3,2) that I need to use to find the value of y''. Can I plug in those values as soon as I substitute the y's?

zepdrix (zepdrix):

Yah that would be a good idea. Plug those values in as soon as possible. You should be able to do some adding and subtracting making all the expansions much easier.

OpenStudy (anonymous):

Well.. when I substitute in the values, I get 0/49 am I wrong, or does this mean there is a local max/min?

zepdrix (zepdrix):

No it doesn't mean something like that :) But hmm, I'm coming up with y''(3,2) = -14/49

zepdrix (zepdrix):

I found it easier to calculate the `y'` on the side, it works out to y'(3,2) = 0, yes?

zepdrix (zepdrix):

\[\large y''(3,2)\quad=\quad \frac{(3\cdot0-2)(8\cdot2-3\cdot3)-(3\cdot2-2\cdot3)(8\cdot0-3)}{(8\cdot2-3\cdot3)^2}\]

zepdrix (zepdrix):

Plugging in 0's for y's, and the rest of the values, I get something like that ^

zepdrix (zepdrix):

That first term in the numerator should give you something besides zero :o

OpenStudy (anonymous):

Doh! I left out the -2 in 3(y')-2 Which would have given me -14/49

OpenStudy (anonymous):

Ok. The question asks if the curve has a local max, min, or neither at the point. How do I determine that?

zepdrix (zepdrix):

Ummmmmmmmmmmmmmm trying to remember...

zepdrix (zepdrix):

There's a formula for second derivative test, i'm trying to remember what it looks like.

OpenStudy (anonymous):

While you think, I'm just going to throw this out there. When I used the calculated y'(8y-3x) = -3 (and similarly for y'(3y-2x) via wolfram alpha and then evaluated, I came up with the same answer. Is that just coincidence?

zepdrix (zepdrix):

I'm not sure what you calculated there. Can you paste the link from wolfram?

OpenStudy (anonymous):

Also, it's similarly used at http://www.derivative-calculator.net/#expr=%283y-2x%29%2F%288y-3x%29&showsteps=1

OpenStudy (anonymous):

I'm referring to earlier when I asked, "Ok, just a quick quesiton. When I plugged this into an online calculator *shame shame* it gave me that (3y-2x)'=-2. Is that true?"

OpenStudy (anonymous):

Sorry if I'm getting confusing here. I'm a little tired ;)

zepdrix (zepdrix):

Mmm ok I'm trying to see what's going on here..

zepdrix (zepdrix):

Ya I can't figure out why all these sites are calculating the partial derivative :( hmm have you learned about partial derivatives at this point? :o It's something you learn about in multi-variable calculus

OpenStudy (anonymous):

Not that I know of! That's why I was asking.. I was afraid I may have dozed off when a shortcut was thought!

OpenStudy (anonymous):

taught* rather

zepdrix (zepdrix):

bah i dunno :3 now you're making me tired lol

OpenStudy (anonymous):

Well, it's 1 a.m. you have the right to be tired! But any last thoughts on the local min, max, or neither bit? If not I'll just leave it for the first couple of minutes when I get to class :P

zepdrix (zepdrix):

does the point have a max, min or neither, I guess since y'(3,2) = 0, we can determine that it IS a critical point and may be a max or min.

zepdrix (zepdrix):

And Ummmm let's see..

zepdrix (zepdrix):

There is a second derivative test, which would tell us if the function is concave up or down at the given point. ~If it's concave up, then the point is a minimum) ~If it's concave down, the point is a maximum) ~I guess it could also be neither, just a critical point that an inflection point. hmm,

OpenStudy (anonymous):

Concavity Theorem: If the function f is twice differentiable at x=c, then the graph of f is concave upward at (cf(c)) if f(c)0 and concave downward if f(c)0.

zepdrix (zepdrix):

Woops your inequality signs didn't show up when pasting :( lol

OpenStudy (anonymous):

Does this mean that since y"=-17/49 it's concavity is down?

zepdrix (zepdrix):

Yes that's what it should mean! :)

OpenStudy (anonymous):

So it's a local max

zepdrix (zepdrix):

Ya that sounds right.. i think! :) lol

OpenStudy (anonymous):

Sounds good to me! Thanks a ton for the help!

zepdrix (zepdrix):

np \c:/